1. A man is standing on the edge of a 20.0 m high cliff. He throws a rock vertically with an initial velocity of

10.0 m/s. How long does it take to reach its max height?

2. A coin is dropped in a vacuum tube, find its displacement after 0.30 seconds.

3. An object dropped on planet X falls 64 m in 4 seconds. Find the acceleration due to gravity on planet X.

4. A basketball player jumped straight up to grab a rebound. If she was in the air for 0.80 seconds, how

high did he jump?

5. A ball drops from a tall building. The ball strikes the sidewalk 2.0 seconds later. What is the velocity of

the ball when it hits the sidewalk?

6. A stone is thrown vertically upward with a speed of 12.0 m/s from the edge of a cliff 70.0 m high. How

long does it take to reach the bottom of the cliff?

7. An object initially at rest falls freely near Earth's surface. How long does it take the object to attain a

speed of 200 m/s?

8. What is the velocity of an object that has been falling freely in a vacuum for 2 seconds?


1
Expert's answer
2021-11-01T12:54:27-0400

"1."

"v_0= 10\frac{m}{s}"

"h_0= 20m"

"g\approx10\frac{m}{s^2}"

"t_{h_{max}}= \frac{v_0}{g}= \frac{10}{10}=1s"

"\text{Answer: }1s"

"2."

"t = 0.3s"

"g\approx10\frac{m}{s^2}"

"v_0=0"

"h_0= 0"

"h_1 = v_0t-\frac{gt^2}{2}= -\frac{10*0.3^2}{2}=-0,45m"

"s = h_0-h_1= 0.45m"

"\text{Answer: }0.45m"

"3."

"h_0 =0"

"t = 4s"

"v_0=0"

"h_1= -64m"

"h = v_0t - \frac{at^2}{2}"

"a = -\frac{2h}{t^2}=\frac{128}{16}=8\frac{m}{s^2}"

"\text{Answer: }8\frac{m}{s^2}"

"4."

"h_0= 0"

"t =0.8s"

"g\approx10\frac{m}{s^2}"

"t_{up}=\frac{t}{2}=0.4s"

"h_{max}= \frac{gt_{up}^2}{2}=\frac{10*0.16}{2}=0.8m"

"\text{Answer: }0.8m"

"5."

"g\approx10\frac{m}{s^2}"

"v_0 =0"

"t = 2s"

"v = v_0+gt = 20\frac{m}{s}"

"\text{Answer: }20\frac{m}{s}"

"6."

"h_0 = 70m"

"v_0= 12\frac{m}{s}"

"g\approx10\frac{m}{s^2}"

"h_1=0"

"h_1= h_0+v_0t-\frac{gt^2}{2}"

"5t^2-12t-70 = 0"

"t \approx5.13s"

"\text{Answer: }5.13s"

"7."

"g\approx10\frac{m}{s^2}"

"v_0= 0"

"v_1=200\frac{m}{s}"

"v_1 = v_0 +gt"

"t =\frac{v_1}{g}=\frac{200}{10}=20s"

"\text{Answer: }20s"

"8."

"v_0= 0"

"t=2s"

"g\approx10\frac{m}{s^2}"

"v = v_0+gt=20\frac{m}{s}"

"\text{Answer: }20\frac{m}{s}"